# Multi-Agent Guided Policy Optimization

### v1

@@ Abstract
Due to practical constraints such as partial observability and limited communication, Centralized Training with Decentralized Execution (CTDE) has become the dominant paradigm in cooperative Multi-Agent Reinforcement Learning (MARL). However, existing CTDE methods often underutilize centralized training or lack theoretical guarantees. We propose Multi-Agent Guided Policy Optimization (MAGPO), a novel framework that better leverages centralized training by integrating centralized guidance with decentralized execution. MAGPO uses an auto-regressive joint policy for scalable, coordinated exploration and explicitly aligns it with decentralized policies to ensure deployability under partial observability. We provide theoretical guarantees of monotonic policy improvement and empirically evaluate MAGPO on 43 tasks across 6 diverse environments. Results show that MAGPO consistently outperforms strong CTDE baselines and matches or surpasses fully centralized approaches, offering a principled and practical solution for decentralized multi-agent learning. Our code and experimental data can be found in https://github.com/liyheng/MAGPO.
@@ 1 Introduction
Cooperative Multi-Agent Reinforcement Learning (MARL) provides a powerful framework for solving complex real-world problems such as autonomous driving [53], traffic management [34], and robot swarm coordination [16, 49]. However, MARL faces two fundamental challenges: the exponential growth of the joint action space with the number of agents, which hinders scalability, and the requirement for decentralized execution under partial observability, which complicates policy learning.
@@
A widely adopted solution is Centralized Training with Decentralized Execution (CTDE) [26, 18], where agents are trained using privileged global information but execute independently based on local observations. CTDE forms the foundation of many state-of-the-art MARL algorithms and typically incorporates a centralized value function to guide decentralized policies or utility functions during training. This setup allows algorithms to benefit from global context without violating the constraints of decentralized deployment.
Parallel efforts in single-agent RL have explored similar ideas in the context of Partially Observable Markov Decision Processes (POMDPs) [25]. Two main approaches have emerged: asymmetric actor-critic [29], which uses full-state information in the critic but restricts the actor to partial observations; and teacher-student learning, where a teacher policy trained with privileged information supervises a student policy that learns to act under partial observability.
These insights have recently inspired the Centralized Teacher with Decentralized Student (CTDS) paradigm in MARL [51]. CTDS combines a centralized critic with a teacher policy that outputs joint actions based on the global state. These actions are then distilled into decentralized policies for execution. While CTDS shows promise—particularly in enabling coordinated exploration and better utilization of centralized training—it faces two key challenges. First, training a centralized teacher remains difficult due to the exponential size of the joint action space. Recent work proposes auto-regressive joint policies [44, 23], where agents act sequentially conditioned on previous actions, mitigating this issue. Second, even with a strong teacher, the decentralized policies may suffer from the imitation gap [43]: student policies operating under partial observability may fail to replicate the teacher’s behavior, leading to degraded performance. In MARL, this issue is exacerbated by policy asymmetry, where the space of decentralized behaviors cannot fully capture the teacher’s joint strategy.
@@
To overcome these limitations, we propose Multi-Agent Guided Policy Optimization (MAGPO), a novel framework that bridges centralized training and decentralized execution through a principled and MARL-specific design. MAGPO explicitly addresses the policy asymmetry problem—unique to multi-agent settings—by constraining a centralized, auto-regressive guider policy to remain closely aligned with decentralized learners throughout training. This alignment ensures that the coordination strategies developed under centralized supervision remain realizable by decentralized policies, thus mitigating the imitation gap that undermines prior CTDS approaches. Unlike a direct extension of single-agent GPO [20], MAGPO introduces structural mechanisms tailored to multi-agent learning, including sequential joint action modeling and decentralized-aligned updates, while preserving scalability and parallelism. We provide theoretical guarantees of monotonic policy improvement and empirically evaluate MAGPO across 43 tasks in 6 diverse environments. Results show that MAGPO consistently outperforms strong CTDE baselines and even matches or exceeds fully centralized methods, establishing it as a theoretically grounded and practically deployable solution for MARL under partial observability.
@@ 2.1 Formulation
We consider Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [25] in modeling cooperative multi-agent tasks. The Dec-POMDP is characterized by the tuple $\langle\mathcal{N},\mathcal{S},\mathcal{A},r,\mathcal{P},\mathcal{O},\mathcal{Z},\gamma\rangle$, where $\mathcal{N}$ is the set of agents, $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $r$ is the reward function, $\mathcal{P}$ is the transition probability function, $\mathcal{Z}$ is the individual partial observation generated by the observation function $\mathcal{O}$, and $\gamma$ is the discount factor. At each timestep, each agent $i\in\mathcal{N}$ receives a partial observation $o_{i}\in\mathcal{Z}$ according to $\mathcal{O}(s;i)$ at state $s\in\mathcal{S}$. Then, each agent selects an action $a_{i}\in\mathcal{A}$ according to its action-observation history $\tau_{i}\in(\mathcal{Z}\times\mathcal{A})^{*}$, collectively forming a joint action denoted as $\bm{a}$. The state $s$ undergoes a transition to the next state $s^{\prime}$ in accordance with $\mathcal{P}(s^{\prime}|s,\bm{a})$, and agents receive a shared reward $r$. Assuming an initial state distribution $\rho\in\Delta(\mathcal{S})$, the goal is to find a decentralized policy $\bm{\pi}=\{\pi_{i}\}_{i=1}^{n}$ that maximizes the expected cumulative return:
$V_{\rho}(\pi)\triangleq\mathbb{E}_{s\sim\rho}[V_{\bm{\pi}}(s)]=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{o}\sim\rho].$ (1)
This work follows the Centralized Training with Decentralized Execution (CTDE) paradigm [26, 18]. During training, CTDE allows access to global state to stabilize learning. However, during execution, each agent operates independently, relying solely on its local action-observation history.
@@ 2.2 Related Works
CTDE methods can be broadly categorized into value-based and policy-based approaches. Value-based methods typically employ a joint value function conditioned on the global state and joint action, alongside individual utility functions based on local observations and actions. These functions often satisfy the Individual-Global-Max (IGM) principle [35], ensuring that the optimal joint policy decomposes into locally optimal policies. This line of work is known as value factorization, and includes methods such as VDN [36], QMIX [30], QTRAN [35], QPLEX [40], and QATTEN [46]. Policy-based methods, in contrast, typically use centralized value functions to guide decentralized policies, allowing for direct extensions of single-agent policy gradient methods to multi-agent settings. Notable examples include COMA [11], MADDPG [22], MAA2C [27] and MAPPO [47]. Additionally, hybrid methods that combine value factorization with policy-based training have been proposed, such as DOP [41], FOP [50], and FACMAC [28]. While CTDE has achieved strong empirical performance, most existing methods leverage global information only through the value function. We refer to these as vanilla CTDE methods, as they do not fully exploit the potential of centralized training.
@@
More recently, researchers have explored extending the teacher-student framework from single-agent settings to multi-agent systems, leading to the Centralized Teacher with Decentralized Students (CTDS) paradigm [51, 5, 54]. In this framework, a centralized teacher policy—accessing global state and acting jointly—collects high-quality trajectories and facilitates more coordinated exploration. CTDS methods offer stronger supervision than vanilla CTDE methods. However, due to observation asymmetry [43] and policy space mismatch, the learned decentralized policies may still suffer from suboptimal performance—issues that we explore further in the next section.
@@
In contrast to vanilla CTDE and CTDS methods—many of which lack theoretical guarantees—another line of research focuses on Heterogeneous Agent Reinforcement Learning (HARL), where agents are updated sequentially during training [52]. This formulation underpins algorithms such as HATRPO and HAPPO [19] and HASAC [21]. While HARL provides better theoretical guarantees and stability, it requires agents to be heterogeneous and updated one at a time. As a result, these methods lack parallelism which is important in large-scale MARL tasks and cannot exploit parameter sharing, which has proven effective in many MARL applications [14, 37, 6].
@@
Centralized Training with Centralized Execution (CTCE) approaches treat the multi-agent system as a single-agent problem with a combinatorially large action space. Beyond directly applying single-agent RL algorithms to MARL, a promising direction in CTCE has been to use transformers [39] to frame multi-agent trajectories as sequences [4]. This has led to the development of powerful transformer-based methods such as Updet [15], Transfqmix [12], and other offline methods [24, 38, 48]. Two representative online methods are Multi-Agent Transformer (MAT) [44] and Sable [23], which currently achieve state-of-the-art performance in cooperative MARL tasks. CTCE methods offer strong theoretical guarantees [44] and impressive empirical results. However, they fall short in practical settings that demand decentralized execution, where each agent must act based solely on its local observation and policy.
@@
3 Problems of CTDS
The first issue—asymmetric observation space—is shared with single-agent POMDPs involving privileged information and has been widely studied in prior work [42, 43, 33, 20]. We only briefly outline it here. When the teacher policy relies on privileged information unavailable (and uninferable) to the student, the student cannot faithfully imitate the teacher’s behavior. In this case, the student tends to learn the statistical average of the teacher’s actions conditioned on the observable input $o$ [43, 42]. This averaged behavior can be significantly suboptimal.
The second issue—asymmetric policy space—is unique to the multi-agent setting. It stems from the structural difference between the teacher’s policy (typically joint and expressive) and the students’ policies (factorized and decentralized). We illustrate this challenge through a simple didactic example in Figure 1. Consider a cooperative task where three agents must each output an integer from 0 to 4, such that their sum equals a target value—here, 10. Each agent acts once, and the system succeeds only if the combined sum is exactly 10. We compare three MARL frameworks:
@@ 3 Problems of CTDS
(A) Vanilla CTDE. In this setting, agents share a centralized value function but act independently using decentralized policies. Suppose all three agents use the same policy: $\pi^{i}(\cdot|10)=[0,0,0,1,0]$, meaning each agent outputs the number 3. The resulting sum is 9, which is incorrect. As each agent observes the same global state and optimizes the same objective, they may all increase their action to 4 in the next update, resulting in 12, still failing to meet the target. Since the setting is fully symmetric and lacks inter-agent coordination signals, agents struggle to resolve who should adjust their actions. This leads to classic miscoordination, where agents must rely on trial-and-error and memorize the rare successful configuration to coordinate effectively.
@@
(B) CTCE. Here, agents act sequentially, and each observes the previous agents’ actions before choosing their own. Suppose the first agent updates its action to 4. The second agent, seeing this, picks 3. The third agent observes both previous actions (4 and 3) and computes the correct final action as 3. The task is completed successfully. Sequential execution turns the multi-agent coordination problem into a single-agent decision-making process over a joint policy. Coordination becomes easy and stable, requiring no repeated attempts. The drawback is that this setting assumes centralized execution, which is infeasible in many real-world applications that demand decentralization.
(C) CTDS. Now suppose we take a successful CTCE policy from (B) and distill it into decentralized student policies. If the teacher policy is deterministic and easily factorizable (e.g., agents always output [4,3,3]), CTDS learns an optimal decentralized solution. However, due to the flexibility of the CTCE policy, it might use a stochastic joint strategy. For instance, the first agent samples action $x$ from $\pi_{1}(\cdot|10)=[0,0,0,0.5,0.5]$, randomly picking 3 or 4. The second agent always picks 3. The third agent, having observed both prior actions, computes the final value as $7-x$, ensuring the total always equals 10. While this policy is optimal for CTCE, it is not factorizable into independent agent policies. If CTDS directly imitates this joint behavior, it may yield:
$\pi_{1}(\cdot|10)=[0,0,0,0.5,0.5],\ \pi_{2}(\cdot|10)=[0,0,0,1,0],\ \pi_{3}(\cdot|10)=[0,0,0,0.5,0.5],$
leading to cases like [4,3,4], summing to 11—thus only achieving 50% success. This highlights the core failure mode: joint coordination encoded in the teacher policy is lost when forced into decentralized policies.
This example illustrates that although CTDS can outperform vanilla CTDE methods such as MAPPO and MAA2C, it remains vulnerable when the teacher policy lies outside the expressiveness of the decentralized policy class—even when observation asymmetry is not a concern. To address this, we propose a new approach that constrains the teacher’s policy during training, preventing it from exploiting unrepresentable coordination strategies, while still enabling it to guide decentralized learners effectively. We introduce this method in the next section.

### v2

@@ Abstract
Due to practical constraints such as partial observability and limited communication, Centralized Training with Decentralized Execution (CTDE) has become the dominant paradigm in cooperative Multi-Agent Reinforcement Learning (MARL). However, existing CTDE methods often underutilize centralized training or lack theoretical guarantees. We propose Multi-Agent Guided Policy Optimization (MAGPO), a novel framework that better leverages centralized training by integrating centralized guidance with decentralized execution. MAGPO uses an autoregressive joint policy for scalable, coordinated exploration and explicitly aligns it with decentralized policies to ensure deployability under partial observability. We provide theoretical guarantees of monotonic policy improvement and empirically evaluate MAGPO on 43 tasks across 6 diverse environments. Results show that MAGPO consistently outperforms strong CTDE baselines and matches or surpasses fully centralized approaches, offering a principled and practical solution for decentralized multi-agent learning. Our code and experimental data can be found in https://github.com/liyheng/MAGPO.
@@ 1 Introduction
Cooperative Multi-Agent Reinforcement Learning (MARL) provides a powerful framework for solving complex real-world problems such as autonomous driving (Zhou et al., 2020), traffic management (Singh et al., 2020), and robot swarm coordination (Hüttenrauch et al., 2019; Zhang et al., 2021a). However, MARL faces two fundamental challenges: the exponential growth of the joint action space with the number of agents, which hinders scalability, and the requirement for decentralized execution under partial observability, which complicates policy learning.
@@
A widely adopted solution is Centralized Training with Decentralized Execution (CTDE) (Oliehoek et al., 2008; Kraemer & Banerjee, 2016), where agents are trained using privileged global information but execute independently based on local observations. CTDE forms the foundation of many state-of-the-art MARL algorithms and typically incorporates a centralized value function to guide decentralized policies or utility functions during training. This setup allows algorithms to benefit from global context without violating the constraints of decentralized deployment.
Parallel developments in single-agent RL have explored analogous ideas under the Partially Observable Markov Decision Process (POMDP) setting (Oliehoek & Amato, 2016). Two representative paradigms have emerged. The first is asymmetric actor–critic (Pinto et al., 2018), where the critic has access to full state information during training while the actor is restricted to partial observations. Conceptually, this setup can be viewed as the single-agent counterpart of CTDE: centralized information is exploited in training but not used at execution time. The second paradigm is teacher–student learning, where a teacher policy trained with privileged information provides direct behavioral supervision to a student policy that must operate under partial observability. Unlike asymmetric actor–critic, which transfers value information, teacher–student methods transfer action-level knowledge, potentially enabling stronger guidance.
While asymmetric designs have long been embedded in CTDE-style MARL algorithms through centralized critics, the teacher–student paradigm has only recently been extended to multi-agent settings. This has led to the Centralized Teacher with Decentralized Student (CTDS) framework (Zhao et al., 2024), which augments CTDE with an explicit centralized teacher policy. In CTDS, a teacher outputs joint actions based on the global state, and decentralized student policies are trained to imitate these actions.
Although CTDS holds the promise of leveraging centralized coordination more effectively than value-based CTDE methods, it introduces structural challenges that are particularly pronounced in MARL. First, learning a centralized teacher over the joint action space suffers from poor scalability, as the space grows exponentially with the number of agents. Second, even if a strong teacher is obtained, decentralized students may suffer from a fundamental imitation gap (Weihs et al., 2021). In multi-agent settings, this gap is exacerbated by policy asymmetry: the centralized teacher conditions on global state and joint context, whereas decentralized students must act based solely on local observations. As a result, the space of realizable decentralized joint behaviors may not contain the teacher’s strategy, leading to unavoidable performance degradation.
@@
To overcome these limitations, we propose Multi-Agent Guided Policy Optimization (MAGPO), a novel framework that bridges centralized training and decentralized execution through a principled and MARL-specific design. MAGPO addresses the scalability and policy asymmetry problem by constraining a centralized, autoregressive guider policy to remain closely aligned with decentralized learners throughout training. The guider policy allows agents to act sequentially conditioned on previous actions, utilizing global information and coordinated data collection (Wen et al., 2022; Mahjoub et al., 2025). The alignment ensures that the coordination strategies developed under centralized supervision remain realizable by decentralized policies, thus mitigating the imitation gap that undermines prior CTDS approaches. Unlike a direct extension of single-agent GPO (Li et al., 2025), MAGPO introduces structural mechanisms tailored to multi-agent learning, including sequential joint action modeling and decentralization-aligned updates, while preserving scalability and parallelism. We provide theoretical guarantees of monotonic policy improvement and empirically evaluate MAGPO across 43 tasks in 6 diverse environments. Results show that MAGPO consistently outperforms strong CTDE baselines and even matches or exceeds fully centralized methods, establishing it as a theoretically grounded and practically deployable solution for MARL under partial observability.
@@ 2.1 Formulation
We consider Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek & Amato, 2016) in modeling cooperative multi-agent tasks. The Dec-POMDP is characterized by the tuple $\langle\mathcal{N},\mathcal{S},\mathcal{A},r,\mathcal{P},\mathcal{O},\mathcal{Z},\gamma\rangle$, where $\mathcal{N}$ is the set of agents, $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $r$ is the reward function, $\mathcal{P}$ is the transition probability function, $\mathcal{Z}$ is the individual partial observation generated by the observation function $\mathcal{O}$, and $\gamma$ is the discount factor. At each timestep, each agent $i\in\mathcal{N}$ receives a partial observation $o_{i}\in\mathcal{Z}$ according to $\mathcal{O}(s;i)$ at state $s\in\mathcal{S}$. Then, each agent selects an action $a_{i}\in\mathcal{A}$ according to its action-observation history $\tau_{i}\in(\mathcal{Z}\times\mathcal{A})^{*}$, collectively forming a joint action denoted as $\bm{a}$. The state $s$ undergoes a transition to the next state $s^{\prime}$ in accordance with $\mathcal{P}(s^{\prime}|s,\bm{a})$, and agents receive a shared reward $r$. Assuming an initial state distribution $\rho\in\Delta(\mathcal{S})$, the goal is to find a decentralized policy $\bm{\pi}=\{\pi_{i}\}_{i=1}^{n}$ that maximizes the expected cumulative return:
$V_{\rho}(\pi)\triangleq\mathbb{E}_{s\sim\rho}[V_{\bm{\pi}}(s)]=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}\sim\rho].$ (1)
This work follows the Centralized Training with Decentralized Execution (CTDE) paradigm (Oliehoek et al., 2008; Kraemer & Banerjee, 2016). During training, CTDE allows access to global state to stabilize learning. However, during execution, each agent operates independently, relying solely on its local action-observation history.
In centralized training, we can therefore optimize a joint policy $\bm{\mu}(\bm{a}|s)$ that coordinates all agents while enabling joint exploration. To update this joint policy, we will consider the policy mirror descent (PMD) objective (Shani et al., 2019):
$\mu^{(k+1)}=\underset{\mu}{\arg\max}\left\{\eta_{k}\langle\nabla V_{\rho}(\mu^{(k)}),\mu\rangle-\frac{1}{1-\gamma}D_{d_{\rho}(\mu^{(k)})}(\mu,\mu^{(k)})\right\},$ (2)
where $\eta_{k}$ is the step size, $\rho\in\Delta(\mathcal{S})$ is an arbitrary state distribution, $d_{\rho}(\mu^{(k)})$ is the discounted state-visitation distribution under $\mu^{(k)}$, and $D_{d_{\rho}(\mu^{(k)})}$ denotes the corresponding weighted Bregman divergence. Conceptually, PMD can be viewed as a class of preconditioned policy gradient methods. Choosing the Bregman divergence to be the Euclidean distance or the Kullback–Leibler (KL) divergence recovers projected policy gradient and natural policy gradient (NPG) (Kakade, 2001), respectively. Thus, PMD provides a unifying framework that encompasses many modern policy-based RL algorithms, including TRPO (Schulman et al., 2015a) and PPO (Schulman et al., 2017). We build upon this formulation to develop our theoretical results in Section 4.1.
@@ 2.2 Related Works
CTDE methods can be broadly categorized into value-based and policy-based approaches. Value-based methods typically employ a joint value function conditioned on the global state and joint action, alongside individual utility functions based on local observations and actions. These functions often satisfy the Individual-Global-Max (IGM) principle (Son et al., 2019), ensuring that the optimal joint policy decomposes into locally optimal policies. This line of work is known as value factorization, and includes methods such as VDN (Sunehag et al., 2017), QMIX (Rashid et al., 2020), QTRAN (Son et al., 2019), QPLEX (Wang et al., 2021a), and QATTEN (Yang et al., 2020). Policy-based methods, in contrast, typically use centralized value functions to guide decentralized policies, allowing for direct extensions of single-agent policy gradient methods to multi-agent settings. Notable examples include COMA (Foerster et al., 2018), MADDPG (Lowe et al., 2017), MAA2C (Papoudakis et al., 2020) and MAPPO (Yu et al., 2022). Additionally, hybrid methods that combine value factorization with policy-based training have been proposed, such as DOP (Wang et al., 2021b), FOP (Zhang et al., 2021b), and FACMAC (Peng et al., 2021). While CTDE has achieved strong empirical performance, most existing methods leverage global information only through the value function. We refer to these as vanilla CTDE methods, as they do not fully exploit the potential of centralized training.
@@
More recently, researchers have explored extending the teacher-student framework from single-agent settings to multi-agent systems, leading to the Centralized Teacher with Decentralized Students (CTDS) paradigm (Zhao et al., 2024; Chen et al., 2024; Zhou et al., 2025). In this framework, a centralized teacher policy—accessing global state and acting jointly—collects high-quality trajectories and facilitates more coordinated exploration. CTDS methods offer stronger supervision than vanilla CTDE methods. However, due to observation asymmetry (Weihs et al., 2021) and policy space mismatch, the learned decentralized policies may still suffer from suboptimal performance—issues that we explore further in the next section.
@@
In contrast to vanilla CTDE and CTDS methods—many of which lack theoretical guarantees—another line of research focuses on Heterogeneous Agent Reinforcement Learning (HARL), where agents are updated sequentially during training (Zhong et al., 2023). This formulation underpins algorithms such as HATRPO and HAPPO (Kuba et al., 2022) and HASAC (Liu et al., 2025). While HARL provides better theoretical guarantees and stability, it requires agents to be heterogeneous and updated one at a time. As a result, these methods lack parallelism which is important in large-scale MARL tasks and cannot exploit parameter sharing, which has proven effective in many MARL applications (Gupta et al., 2017; Terry et al., 2020; Christianos et al., 2021a).
@@
Centralized Training with Centralized Execution (CTCE) approaches treat the multi-agent system as a single-agent problem with a combinatorially large action space. Beyond directly applying single-agent RL algorithms to MARL, a promising direction in CTCE has been to use transformers (Vaswani et al., 2017) to frame multi-agent trajectories as sequences (Chen et al., 2021). This has led to the development of powerful transformer-based methods such as Updet (Hu et al., 2021), Transfqmix (Gallici et al., 2023), and other offline methods (Meng et al., 2023; Tseng et al., 2022; Zhang et al., 2022). Two representative online methods are Multi-Agent Transformer (MAT) (Wen et al., 2022) and Sable (Mahjoub et al., 2025), which currently achieve state-of-the-art performance in cooperative MARL tasks. CTCE methods offer strong theoretical guarantees (Wen et al., 2022) and impressive empirical results. However, they fall short in practical settings that demand decentralized execution, where each agent must act based solely on its local observation and policy.
3 Problems of Current CTDS
The first challenge—asymmetric observation spaces—is shared with single-agent POMDPs involving privileged information and has been extensively studied in prior work (Warrington et al., 2020; Weihs et al., 2021; Shenfeld et al., 2023; Li et al., 2025). When the teacher relies on privileged information that is unavailable (and not inferable) to the student, the student cannot faithfully reproduce the teacher’s behavior. Instead, it learns to approximate the conditional expectation of the teacher’s action given the observable input (Weihs et al., 2021; Warrington et al., 2020). This typically results in an “averaged” behavior that can be significantly suboptimal.
The second challenge—asymmetric policy spaces—is unique to the multi-agent setting. It arises from the structural mismatch between the teacher’s policy (typically joint and expressive) and the students’ policies (factorized and decentralized). We illustrate this challenge through a simple example shown in Figure 1. Consider a cooperative task where three agents must each output an integer such that their sum equals a target value of $10$. Each agent acts once, and the system succeeds only if the total sum is exactly $10$. We compare three MARL frameworks:
@@ 3 Problems of Current CTDS
(A) Vanilla CTDE. Agents share a centralized value function but act independently via decentralized policies. Suppose all three agents use the same deterministic policy $\pi^{i}(\cdot|10)=3$, producing a total of 99—which fails the task. Because each agent observes the same global state and optimizes the same objective, they may all simultaneously increase their action to $4$ in the next update, producing $12$, still failing. Lacking inter-agent coordination signals, the agents struggle to determine which one should adjust its action. This leads to classic miscoordination, requiring random trial-and-error exploration and memorization of rare successful configurations to eventually coordinate.
@@
(B) CTCE. Agents act sequentially, observing previous agents’ actions before choosing their own. Suppose the first agent updates its action to $4$ and the second still picks $3$. The third agent, having observed both previous actions, selects $3$ and achieves the correct total. Sequential execution effectively transforms the multi-agent coordination problem into a single-agent decision-making process over a joint policy. This makes coordination straightforward and stable, without repeated random exploration. However, this setting assumes centralized execution, which is often infeasible in real-world applications requiring decentralization.
(C) CTDS. Now consider distilling a successful CTCE policy from (B) into decentralized student policies. If the teacher’s policy is deterministic and factorizable (e.g., always producing $[4,3,3]$), CTDS can recover an optimal decentralized solution. However, a CTCE teacher may exploit stochastic strategies. For instance, the first agent samples $x\in\{3,4\}$ at random, the second agent always outputs $3$, and the third agent—having seen the first two actions—outputs 7x7-x, ensuring the total is always $10$. While this is optimal under CTCE, it cannot be factored into independent policies: CTDS would learn two independent stochastic policies for the first and third agents, leading to failures such as $[4,3,4]$. If the first CTCE agent selects $3$ or $4$ with equal probability, the distilled decentralized policy succeeds only $50\%$ of the time.
This example highlights the core failure mode: coordination patterns encoded in the teacher’s joint policy are lost when forced into a decentralized representation. To address these limitations, we propose a new approach that constrains the teacher’s policy during training, preventing it from exploiting unrepresentable coordination strategies while still allowing it to guide decentralized learners effectively.